OpenStudy (anonymous):

SOLVE THE EQUATION. BE SURE TO SHOW YOUR WORK. ALL FINAL ANSWERS MUST BE IN SIMPLEST FORM. N-1/6=8

OpenStudy (anonymous):

The equation is?

OpenStudy (anonymous):

n - 1/6=8

OpenStudy (anonymous):

add 1/6 to both sides of the equation

OpenStudy (anonymous):

\[n = 8 - \frac{1}{6}=\frac{47}{6}\]

OpenStudy (anonymous):

\[n = 8+\frac{1}{6} = \frac{49}{6}\]

OpenStudy (anonymous):

sorry hehe

OpenStudy (anonymous):

Check our answer by plugging 49/6 back into the original equation

OpenStudy (anonymous):

ok thanks